Please show all the work! Thank you so much!
The following is the Shareholders Equity section of the BillyBoB Corporation at December 31, 2017: Preferred Stock, 10% annual dividend, $40 par, 100,000 shares authorized, 7,000 shares issued& outstanding280,000 Common Stock, $5 par, 750,000 shares authorized, 112,000 issued & outstanding.. Paid-In Capital in excess of par on preferred...70,000 Paid-In Capital in excess of par on common210,000 560,000 Retained Earnings 1,280,000 January 20, 2018 BillyBoB Corp issued 38,000 shares of common stock for $456,000 cash. February 5, 2018 The price of common stock has increased 25 % since the January 20, 2015 sale. BillyBoB Corp. declared a 10% common stock dividend to shareholders of record February 10th, payable February 18th. BillyBoB Corp buys back 13,200 of its own common shares at an average price of $18 cash per share. February 28, 2018 BillyBoB Corp. declared a cash dividend on common stock of 25 cents per share to shareholders of record March 10th, payable March 28th. In addition, preferred dividends are declared and paid for the quarter. March 4, 2018 Required: 1. 2. Journalize the above transactions omitting explanations. Prepare a Shareholders Equity section of the BillyBoB Corporation at March 31, 2018, assuming the net income for January, February, and March is $82,000.
Answer #1
1) Journal Entries:
Date Accounts title Debit $ Credit $
20-Jan-18 Cash 456000
Common Stock 190000
Paid in capital in excess of par on CS 266000
5-Feb-18 Stock dividend 225000
Common Stock distributable 75000
Paid in capital in excess of par on CS 150000
(stock dividend declared at MV
(456000/38000)*1.25= 15 and quantity
(112000+38000)*10%=15000 )
18-Feb-18 Common Stock distributable 75000
Common Stock   75000
(being common stock issued)
28-Feb-18 Treasury StocK 237600
Cash 237600
(being buy back of CS 13200@$18)
4-Mar-18 Dividend 37950 (151800*0.25)
Preferred Dividend 7000 (280000*10%*1/4)
Dividend payable 37950
Pref. dividend payable 7000
(165000 - 13800= 151800 CS )
31-Mar-18 Retained Earnings 269950
Stock dividend 225000
Dividend 37950
Preferred Dividend 7000
31-Mar-18 Income Summary 82000
Retained Earnings 82000
2) Shareholders Equity Section as on March 31, 2018 :
Preferred Stock, 10% annual dividend, $40 Par,100000
shares authorised, 7000 shares issued and outstanding $280,000
Common Stock, $5 par, 750000 shares authorised,
165000 shares issued and 151800 outstanding 759000
Paid in capital in excess of par on preferred stock 70000
Paid in capital in excess of par on common stock 626000
Capital contribution $1,735,000
Retained Earnings 1092050
Less: Treasury Stock -237600 854450
Shareholder's Equity = $2,589,450